OpenStudy (henrietepurina):

The equation of a line is shown below. y = - x - 7 What is the equation of a line which is parallel to this line and passes through (1, 79)?

OpenStudy (henrietepurina):

y = -x + 80 y = x - 80 y = 7x - 80 y = -7x + 80

OpenStudy (kc_kennylau):

First of, two lines that are parallel to each other have the same slope

OpenStudy (kc_kennylau):

What's the slope of the line in the question?

OpenStudy (henrietepurina):

ho did you do slope again x - x^1 = m(y - y^2) ?

OpenStudy (kc_kennylau):

The equation in the question is in slope-intercept form, i.e. y=mx+b where m is the slope and b is the y-intercept

OpenStudy (henrietepurina):

ooh oops

OpenStudy (kc_kennylau):

So what's the slope?

OpenStudy (henrietepurina):

wait whats y, y^1 and x, x^2

OpenStudy (kc_kennylau):

Your equation is not correct, it's y-y_1 = m(x-x_1)

OpenStudy (kc_kennylau):

Or written out: \(y-y_1=m(x-x_1)\)

OpenStudy (henrietepurina):

sorry thats what i meant but still what is y and y_1 and x and x_1

OpenStudy (kc_kennylau):

Where \((x_1,y_1)\) is a point on the line and \(m\) is the slope

OpenStudy (kc_kennylau):

All (x,y) that satisfy the equation will be on the line and all (x,y) on the line will satisfy the equation

OpenStudy (henrietepurina):

y - 79 = m(x - 1)

OpenStudy (henrietepurina):

y - 79 = mx - m y - 79 = x 79 = x + y?

OpenStudy (kc_kennylau):

mx-m is not x

OpenStudy (kc_kennylau):

Actually, NOT using point-slope form will be faster

OpenStudy (henrietepurina):

so what should we use?

OpenStudy (kc_kennylau):

slope-intercept form y=mx+b where m is the slope and b is the y-intercept

OpenStudy (henrietepurina):

y = m79 + 1

OpenStudy (kc_kennylau):

Find the slope and the y-intercept in the equation "y=-x-7" first

OpenStudy (henrietepurina):

how if theres no m

OpenStudy (kc_kennylau):

m is a constant

OpenStudy (kc_kennylau):

In this case m is -1

OpenStudy (henrietepurina):

ooh

OpenStudy (kc_kennylau):

Find the option where m is also -1

OpenStudy (kc_kennylau):

Shortcuts for MC hehe

OpenStudy (henrietepurina):

-1 * x - 7

OpenStudy (kc_kennylau):

If there's no m, m is still 1

OpenStudy (henrietepurina):

x = 6?

OpenStudy (kc_kennylau):

Since they are parallel, you can find the option in which the equation also has m=-1

OpenStudy (henrietepurina):

-x - 7

OpenStudy (henrietepurina):

i really dont get this and gotta move along can you pls tell me the answer

OpenStudy (kc_kennylau):

Looks like you're not getting it, I apologize for confusing you

OpenStudy (kc_kennylau):

For example, in the equation y=2x+5

OpenStudy (kc_kennylau):

The slope is 2 and the y-intercept is 5

OpenStudy (kc_kennylau):

Get this?

OpenStudy (henrietepurina):

ahh so the slope was -1?

OpenStudy (kc_kennylau):

Yep

OpenStudy (kc_kennylau):

So find the option where the slope is also -1 :)

OpenStudy (henrietepurina):

A?

OpenStudy (kc_kennylau):

Exactly :)

OpenStudy (henrietepurina):

omg thanks a lawt

OpenStudy (kc_kennylau):

no prawblem

OpenStudy (henrietepurina):

hehe